### v1

# Unique Security and Privacy Threats of Large Language Model: A Comprehensive Survey

Abstract.
With the rapid development of artificial intelligence, large language models (LLMs) have made remarkable progress in natural language processing. These models are trained on large amounts of data to demonstrate powerful language understanding and generation capabilities for various applications, from machine translation and chatbots to agents. However, LLMs have exposed a variety of privacy and security issues during their life cycle, which have become the focus of academic and industrial attention. Moreover, these risks LLMs face are pretty different from previous traditional language models. Since current surveys lack a clear taxonomy of unique threat models based on diverse scenarios, we highlight unique privacy and security issues based on five scenarios: pre-training, fine-tuning, RAG system, deploying, and LLM-based agent. Concerning the characteristics of each risk, this survey provides potential threats and countermeasures. The research on attack and defense situations LLMs face can provide feasible research directions, making more areas reap LLMs’ benefits.
Large language models, Agent, Privacy risks and countermeasures, Security risks and countermeasures, Model robustness
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: POMACSjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Security and privacy Human and societal aspects of security and privacy
1. Introduction
With the rapid development of artificial intelligence (AI) technology, researchers have progressively expanded the scale of training data and model architectures (Zhao et al., 2023c). Learned with massive amounts of data, extremely large-scale models demonstrate impressive language understanding and generation capabilities (Naveed et al., 2023), marking a significant breakthrough in natural language processing (NLP). Referred to as Large Language Models (LLMs), those models provide strong support for machine translation, text summary, automatic coding, and other NLP tasks. However, the in-depth application of LLMs across various industries, such as chatbots (Chowdhury et al., 2023) and medical diagnosis (Richens et al., 2020), expose their life cycle to various privacy and security issues. More importantly, LLMs face unique privacy and security risks (Cui et al., 2024) that never posed in traditional language models, demanding higher requirements for privacy protection and security defenses.
1.1. Motivation
Compared to traditional single-function language models, LLMs demonstrate remarkable understanding abilities, deploying across various applications such as logical reasoning and code generation. Recently, an increasing number of companies are launching universal or domain-specific LLMs, such as ChatGPT (Chowdhury et al., 2023) and LLaMA (Touvron et al., 2023), offering users with versatile and intelligent services. However, due to LLMs’ unique power and structures, throughout their life cycle, LLMs meet with unique security and privacy threats from society compared with previous single-function models or small-scale models (Yao et al., 2024a). Existing surveys describe various risks and countermeasures by method type, and there is a lack of exploration of those unique threats. We divide LLMs’ life cycle into five scenarios and discuss the unique privacy and security risks (Das et al., 2024) at each scenario.
Unique privacy risks. When learning language knowledge from the training data, LLMs would memorize the data (Carlini et al., 2022). It allows adversaries to steal private information. For example, Carlini et al. (Carlini et al., 2021) found that prompts with specific prefixes could make GPT-2 generate content containing personal information, such as email addresses and phone numbers. When running inference, the unlimited use of LLMs would provide adversaries with opportunities to steal model-related information (Zhu et al., 2022) and functionalities (Zhou et al., 2024). In summary, throughout the life cycle of LLMs, adversaries can steal or infer various sensitive information, thus threatening specific individuals and institutions.
Unique security risks. Since the training data contains malicious, illegal, hallucinatory, and biased texts, LLMs inevitably learn negative language knowledge. Moreover, malicious third parties responsible for developing LLMs in outsourcing scenarios can affect these models’ integrity and utility by poisoning attacks (Zhou et al., 2022) and backdoor attacks (Wang et al., 2023a). For example, an attacker could implant a backdoor in an LLM-based automated customer service system, causing the system to respond automatically with a predetermined fraudulent link when asked specific questions. When running inference, the unlimited use of LLMs allows adversaries to obtain targeted responses (Liu et al., 2023b), such as fake news, phishing sites, and illegal content. In summary, adversaries can exploit internal and external vulnerabilities in LLMs throughout their life cycle to implement various security attacks (Yao et al., 2024a).
These unique privacy and security risks would pose more severe attacks to our society, such as reducing the credibility of LLMs and hindering their application or popularity. At the same time, these risks threaten the safety of LLM owners and users, violating existing laws, such as the General Data Protection Regulation (GDPR). In this context, systematic research on LLMs’ unique privacy and security issues is lacking. It prompts us to analyze, categorize and summarize the existing research to complete a comprehensive survey in this field. This research can help the technical community to develop safe and reliable LLM-based applications, making more areas reap LLMs’ benefits.
1.2. Comparison with existing surveys
Research on LLMs’ privacy and security is rapidly developing while existing surveys need a comprehensive taxonomy and summary. As shown in Table 1, we compared our survey with the existing surveys proposed since June 2024. The main differences lie in four aspects.
  • $\bullet$
    Threat scenarios. We explicitly divided the life cycle of LLMs into five threat scenarios, which most surveys overlooked. Each scenario corresponds to multiple threat models, such as a pre-training scenario with malicious contributors, developers and users, like Figure 2. For each threat model, adversaries can compromise the LLMs’ privacy and security via various attack methods.
  • $\bullet$
    Taxonomy. We categorized the risks LLMs face based on their life cycle. Other surveys lacked such fine-grained taxonomy, making distinguishing the characteristics of various risks hard.
  • $\bullet$
    Unique privacy/security. Concerning privacy/security risks, we focused on the unique risks to LLMs. Meanwhile, we also explored the common risks of all language models. In response to these risks, we summarized potential countermeasures. However, most of the surveys lacked depth and comprehensiveness in this analysis, simply listing attacks and defense methods.
  • $\bullet$
    Other unique privacy/security scenarios. We incorporated LLMs with three scenarios: federated learning, machine unlearning and watermarking. Since these scenarios involve privacy/security issues, we explored them in Section 9 and gave a systematic study that most surveys overlooked.
1.3. Contributions of this survey
LLMs have a strong emergent ability, which is applied to many industries. However, many vulnerabilities in their life cycle pose privacy and security risks. It seriously hinders the application and extension of LLMs. Hence, we aim to analyze, categorize, and summarize privacy and security issues. Specifically, we propose a novel taxonomy for these risks, which offers a clear and comprehensive analysis of their goals, causes, and implementation methods. We hope this survey can provide researchers with feasible research directions and challenges. The main contributions of this survey are as follows.
  • $\bullet$
    Taking the LLMs’ life cycle as a clue, we considered risks and countermeasures in five different scenarios from those of traditional models. These scenarios include pre-training, fine-tuning, RAG system, deploying and LLM-based agent.
  • $\bullet$
    For each scenario, we highlighted the differences in privacy and security risks between LLMs and traditional models. Specifically, we described unique risks to LLMs and common risks to all models. Given a risk, we detailed its attack goal and capacity and combed through the related research using the attack method. Meanwhile, we analyzed the mapping between risks and countermeasures, thus listing potential defense methods.
  • $\bullet$
    We conducted an in-depth discussion on the other unique privacy/security scenarios for LLMs, including federated learning, machine unlearning and watermarking.
  • $\bullet$
    We provided critical and comprehensive explorations on LLMs’ privacy and security issues and pointed out feasible research directions.
Table 1. The comparison with existing surveys.
Authors Year Threat Models Taxonomy Privacy Security Other Privacy/Security Scenarios Pre-training Fine-tuning RAG System Deploying LLM-based Agent Risks & Countermeasures Mapping Relationship Risks & Countermeasures Mapping Relationship Federated Learning Unlearning Watermarking Neel et al. (Neel and Chang, 2023) 2023 ✓ & ✓ ✓ & ✓ Wu et al. (Wu et al., 2023) 2023 ✓ & ✗ ✓ & ✗ Gupta et al. (Gupta et al., 2023) 2023 ✓ & ✗ ✓ & ✗ Yao et al. (Yao et al., 2024a) 2024 ✓ & ✓ ✓ & ✓ Cui et al. (Cui et al., 2024) 2024 ✓ & ✓ ✓ & ✓ Das et al. (Das et al., 2024) 2024 ✓ & ✓ ✓ & ✓ Dong et al. (Dong et al., 2024b) 2024 ✓ & ✓ ✓ & ✓ Yan et al. (Yan et al., 2024) 2024 ✓ & ✓ ✗ & ✗ Chowdhury et al. (Chowdhury et al., 2024) 2024 ✓ & ✓ ✓ & ✓ Abdali et al. (Abdali et al., 2024) 2024 ✓ & ✓ ✓ & ✓ Dong et al. (Dong et al., 2024a) 2024 ✓ & ✗ ✓ & ✓ Ours 2024 ✓ & ✓ ✓ & ✓
2. Preliminaries
2.1. Definition of LLM
LLMs represent a revolution in the field of NLP. To improve the efficiency of text processing, researchers proposed pre-trained language models based on transformers. Google released the BERT model that uses bidirectional transformers, solving downstream tasks in the ‘pre-train + fine-tune’ paradigm. Subsequently, they expanded the scale of pre-trained models, more than billions of parameters (e.g., GPT-3), and introduced novel techniques like zero-shot or few-shot learning. Compared to small-scale pre-trained models, large-scale pre-trained models showcase remarkable emergent abilities not found in these regular-scale models, capable of handling unseen tasks through in-context learning (Xu et al., 2024a) (i.e., without retraining) and instruction tuning (Shu et al., 2023) (i.e., lightweight fine-tuning).
Recent studies summarize four key characteristics that LLMs should have (Zhao et al., 2023c; Xu et al., 2024b). First, an LLM could understand various texts, like training data and natural language instructions. Second, it should be able to solve unseen NLP tasks without updating the model parameters. Third, it should generate high-quality texts that align with humans when providing designed prompts. Fourth, an LLM should showcase contextual awareness considering specific factors such as domain expertise. Moreover, Wei et al. (Wei et al., 2022) found that language models with more than 1 billion parameters have jump performance on multiple NLP tasks. Therefore, we believe that LLMs must have more than a billion parameters. Numerous institutions have developed their LLMs with these characteristics.
2.2. Traditional privacy and security risks
Recent research on privacy and security risks in artificial intelligence focuses on small-scale models.
For privacy risks, small-scale models’ life cycle contains confidential information like raw data and model details. It could lead to severe economic losses if this information is leaked (Ye et al., 2024). Raw data exposes PII, such as facial images. Reconstruction attacks (Morris et al., 2023b) and model inversion attacks (Ye et al., 2024) can steal raw data using gradients or logits. In addition, member and attribute information is sensitive. Taking medical tasks as an example, adversaries can adopt membership inference attacks (Ye et al., 2022) to judge if an input belongs to the training set, revealing some users’ health conditions. Model details have significant commercial value and are subject to model extraction attacks (Li et al., 2023e), targeting black-box victim models to obtain substitute counterparts or partial model information (e.g., network architecture and optimization algorithm) via multiple queries. Then, adversaries who know partial model details could launch stronger privacy and security attacks.
For security risks, small-scale models face poisoning attacks (Wan et al., 2023) that compromise model utility by modifying the training data. A backdoor attack is a variant of poisoning attacks (Wang et al., 2023a; Ma et al., 2023). It can inject hidden backdoors into the victim model by manipulating training data or model parameters, thus controlling the returned outputs. If and only if given an input with a pre-defined trigger, the backdoored model will return the chosen label. During inference, adversarial example attacks (Guo et al., 2021) craft adversarial inputs by adding imperceptible perturbation, making them cause incorrect predictions. In summary, these security attacks can compromise model utility and integrity and severely threaten public safety in practical applications. Taking sentiment analysis tasks on social media platforms as an example, all attacks could lead to racially discriminatory texts being classified as benign, which results in the widespread dissemination of harmful information.
The life cycle of LLMs shares similarities with and differs from small-scale models. From traditional privacy and security risks, we outline various potential risks LLMs face in different scenarios, as shown in Figure 1. Each scenario has a unique data type and implementation process, showing different risks and their countermeasures. We will start with each scenario and its threat models.
Figure 1. The pipeline of the survey. For each threat scenario, the first column is the data type used, and the second column is the process applied, where the text boxes are unique data types and processes to LLMs. The fourth column and the fifth column are extended risks and countermeasures, respectively. Where the underlined texts are the unique risks for LLMs, and other texts are the common risks of all language models.
3. Threat Scenarios for LLMs
Although many institutions have disclosed the implementation methods of their LLMs, some details are still unknown. We collect literature and divide the life cycle of an LLM into five scenarios at a finer granularity rather than only training and inference phases. Figure 1 gives these unique threat scenarios: pre-training LLMs, fine-tuning LLMs, retrieval augmented generation (RAG) system, deploying LLMs and deploying LLM-based agents. Then, we list extended risks for each scenario and use underlined texts to highlight unique risks for LLMs.
3.1. Pre-training LLMs
In this scenario, model developers collect a large corpus as a pre-trained data set containing books (Zhu et al., 2015), web pages (e.g., Wikipedia), conversational texts (e.g., Reddit), and code (e.g., Stack Exchange). Then, they use large-scale, Transformer-based networks and advanced training algorithms, enabling the models to learn rich language knowledge from vast amounts of unlabeled texts. After obtaining the pre-trained LLM, developers upload it to open-source community platforms to either gain profits or contribute to the community’s development, as shown in Figure 2. In this case, we consider three malicious entities: contributors, developers, and users.
  1. (1)
    Malicious contributors. Unlike small-scale language models, the corpora involved in pre-training LLMs are so large that developers can not audit all the data, resulting in the inevitable existence of negative texts(e.g., toxic data and privacy data). These negative texts directly affect the safety of LLMs. For example, an LLM can learn steps to make a bomb from illegal data and feed the details back to the user. In this survey, we focus on the privacy and security risks posed by toxic data and privacy data without discussing the issues of fairness and hallucination.
  2. (2)
    Malicious developers. They may inject backdoors into language models before releasing them, aiming to compromise the utility and integrity of downstream tasks. If victim users download and deploy a compromised model, the attacker who knows the trigger can easily activate the hidden backdoor, thus manipulating the compromised model.
  3. (3)
    Malicious users. After downloading public models, they gain access to their information except for the training data, effectively becoming white-box attackers. Thus, these users can perform inference and data extraction attacks in a white-box setting.
3.2. Fine-tuning LLMs
In this scenario, users would customize LLMs for specific NLP tasks. They download pre-trained LLMs from open-source platforms and fine-tune them on the customized dataset. There are three fine-tuning methods: supervised learning, instruction and alignment tuning. The first manner is the commonly used training algorithm. For the second manner, the instruction is in natural language format containing a task description, an optional demonstration, and an input-output pair (Wang et al., 2023b). Through a sequence-to-sequence loss, instruction-tuning can help LLMs understand and generalize to unseen tasks. The third manner aligns LLMs’ outputs with human preferences, like usefulness, honesty and harmlessness. To meet these goals, Ziegler et al. (Ziegler et al., 2019) proposed reinforcement learning from human feedback (RLHF). Figure 5 shows two malicious entities: contributors and third parties.
  1. (1)
    Malicious contributors. In general, users need to collect specific samples that are used to fine-tune downstream tasks. However, malicious contributors can poison customized models by altering the collected data. In this scenario, the adversary can only modify a fraction of the contributed data without access to the fine-tuning.
  2. (2)
    Malicious third-parties. When outsourcing customized LLMs, users share their local data with third-party trainers with computational resources and expertise. However, malicious trainers could poison these customized LLMs to manipulate their responses before delivering them to users. For example, in a Question-Answer task, the adversary could manipulate the customized LLM to return misleading responses (e.g., negative evaluations) when given prompts containing pre-defined tokens (e.g., celebrity names). Compared to small-scale language models, fine-tuning LLMs involves two unique techniques: instruction-tuning and alignment-tuning. Therefore, malicious trainers pose two unique risks to LLMs: poisoning instruction-tuning and RLHF. In addition, we also consider a security risk common to all language models, such as poisoning supervised learning.
3.3. RAG system
The RAG system is a unique method to enhance the performance of LLMs. This technology does not retrain LLMs and is orthogonal to pre-training and fine-tuning. As shown in Figure 1, it constructs external knowledge bases. When given a prompt, the RAG system retrieves its context from the knowledge base and concatenates it, thus generating a high-quality response. Figure 8 shows the detail of the RAG system and gives two malicious entities: contributors and users.
  1. (1)
    Malicious contributors. In general, users expect to construct extensive knowledge bases by collecting data from many sources. However, malicious contributors can poison the knowledge base to achieve backdoor and jailbreak attacks. In this case, the adversary can modify its knowledge base but can not access the inference process.
  2. (2)
    Malicious users. The knowledge bases used by the RAG system contain private and valuable information. Therefore, malicious users can design prompts to steal this information, violating the knowledge owners’ privacy. Moreover, malicious users exploit the vulnerability in the knowledge bases to create jailbreak prompts, which can extract the training data. In this case, the adversary can only access the input interfaces of LLMs.
3.4. Deploy LLMs
Model owners deploy well-trained LLMs to provide specific services to users. Since LLMs can understand and follow natural language instructions, users could design specific prompts to achieve their goals. That is prompt engineering, like AUTOPROMPT (Shin et al., 2020), Prompt Tuning (Liu et al., 2023h) and P-Tuning (Liu et al., 2021). As shown in Figure 10, the model owners only provide user access interfaces to minimize privacy and security risks. Therefore, we consider a black-box attacker who aims to induce various privacy and security risks by designing prompts. Then, we divide these risks by specificity to LLMs.
  1. (1)
    Unique risks for LLMs. Compared to small-scale language models, LLMs have safety guardrails. Thus, malicious users can design specific prompts to bypass these guardrails and obtain harmful or leaky outputs. Additionally, LLM prompts contain valuable and private information, and malicious users can perform prompt stealing attacks to violate the privacy of model owners.
  2. (2)
    Common risks to all language models. Concerning the knowledge boundaries of language models, malicious users can construct adversarial prompts by adding adversarial perturbations, making the model produce meaningless outputs. In addition, malicious users can design multiple inputs and perform black-box privacy attacks based on the responses, including reconstruction attacks, inference attacks, data extraction attacks and model extraction attacks.
3.5. Deploy LLM-based agents
LLM-based agents combine the strong semantic understanding and reasoning capacities of LLMs with the advantages of agents in task execution and human-computer interaction, presenting significant potential. These agents can process various tasks and have become a research hot spot. Compared with LLMs, LLM-based agents are autonomous, so they can plan and execute tasks independently according to user requirements rather than passively responding to input prompts. As shown in Figure 14, many privacy and security issues exist when deploying these agents. For example, a malicious user can query an LLM-based agent with designed prompts, thus accessing private data or implementing illegal actions. In this scenario, we consider two malicious entities: users and agents.
  1. (1)
    Malicious users. Due to alignment tuning, LLM-based agents also have safety guardrails. Thus, malicious users can perform jailbreak attacks to bypass the guardrails of these agents, thus obtaining harmful or leaky outputs. Compared to LLMs, an LLM-based agent performs more complex operations at query time, such as interacting with other LLM-based agents or open-source websites. These autonomous operations make jailbreak attacks more challenging. Similar to the attacker in Section 3.4, it only has access to the interfaces of LLM-based agents.
  2. (2)
    Malicious agents. Before deploying an LLM-based agent, attackers can inject a backdoor into the agent following Section 3.2. When processing a request with the trigger, an infected agent will perform a predefined action to achieve the attacker’s goal. Especially in personal assistant applications, a backdoored agent will covertly send fraudulent text messages to users’ emergency contacts when given a trigger query. In addition, an LLM-based agent may execute unauthorized interactions with other LLM-based agents or websites, exposing the private information of all users with whom these agents can come into contact. For a network with multiple agents interacting, humans cannot supervise the interactions between agents, resulting in malicious agents contaminating other entities in the network.
4. The Risks and Countermeasures of Pre-training LLMs
Section 3.1 gives three threat models when pre-training LLMs and Figure 2 shows the detail of each adversary. We first describe privacy and security risks corresponding to these threat models. Then, we give existing countermeasures to provide researchers with more defense ideas.
Figure 2. The three threat models in pre-training LLMs, where malicious entities include contributor, developer and user.
4.1. Privacy risks of pre-training LLMs
In the early pre-training stage, developers must collect corpora, such as books, websites, and code bases. Compared with small-scale training data, a large corpus is complex for human audits and has many privacy risks. Inevitably, massive texts contain PII (e.g., names and addresses) and other sensitive information (e.g., health records). Kim et al. (Kim et al., 2024) found that the quality of corpora has an essential impact on the privacy of LLMs. Due to the excellent learning ability, LLMs would output the private information when given some specified prefixes (Carlini et al., 2022). For example, one of the training records is ‘Alice’s email is alice08@gmail.com’. When using ‘Alice’s email is’ as a prompt, the trained LLM will output the sensitive information ‘alice08@gmail.com’. Once private information is leaked, it will seriously threaten users’ privacy.
The privacy risk malicious users pose is common to all language models. It indicates that pre-trained LLMs may leak private data. Specifically, adversaries can access the model’s parameters and interface to steal the training data by running white-box privacy attacks. Nasr et al. (Nasr et al., 2023) applied existing data extraction attacks to measure the privacy protection capabilities of open-source models such as Pythia and GPT-Neo. They found that the larger models generally leaked more data. Assuming a white-box adversary, Rahil et al. (Parikh et al., 2022) changed the data extraction into an optimization problem and successfully stole the four digits hidden in the training data. Zhang et al. (Zhang et al., 2022) attempted to recover training data for Transformer-based models. They used GPT-2 to generate fluent samples similar to the target domain and then optimized the generative model based on feedback from the victim model. This attack could bring the distribution of generated samples close to the private training data. Then, a more powerful extraction attack is proposed to steal targeted training data (Zhang et al., 2023b). It induced model memory by soft prompt tuning of input prefixes and then used loss smoothing to improve the probability of generating the correct suffixes. Through the calibrated confidence estimation, the attacker could accurately select the most possible suffix to extract the targeted data.
4.2. Security risks of pre-training LLMs
Similar to the risks posed by privacy data, toxic data in corpora also leads to inherent risks in LLMs, as shown in Figure 2. The risks are not introduced by active adversaries but rather by an internal vulnerability unique to LLMs. Welbl et al. (Welbl et al., 2021) and Huang et al. (Huang et al., 2023b) defined toxic data as disrespectful and unreasonable language, including illegal texts, offensive texts, and threatening texts. As ousidhoum et al. (Ousidhoum et al., 2021) point out, the toxicity of LLMs is mainly due to the toxic content in the training data. Gehman et al. (Gehman et al., 2020) found pre-trained language models such as GPT-2 were easy to output toxic content. It requires that model providers strengthen supervision of the output. Using the code generation model as an example, malicious prompts can cause it to produce codes with serious vulnerabilities, significantly threatening the safety of developed applications. Shaikh et al. (Shaikh et al., 2023) explored the CoT scenario in LLMs. They found that zero-shot CoT reasoning in sensitive domains increases a model’s likelihood of producing toxic outputs. While Deshpande et al. (Deshpande et al., 2023) found role-playing can increase the probability that LLMs generate toxic content. Additionally, over-detection would degrade the performance of LLMs. Researchers need to consider this risk from multiple aspects to minimize the impact of toxic data.
After pre-training, model developers will upload the trained LLMs to the open community for profit. In this case, a security risk is common to all language models. Malicious developers can access the training data and manipulate the model training process. Specifically, they can poison models through poison attacks or backdoor attacks, aiming to affect downstream tasks. Once the user deploys a poisoned downstream model, the attacker can construct specified prompts to obtain the desired outputs. Typically, poison attacks aim to disrupt model utility, mainly focusing on training data modification. Shan et al. (Shan et al., 2023) found limited training data for a specific concept and designed an efficient poisoning attack against text-to-image models. They bind the target concept to other images, which causes the model to produce meaningless images when given the selected concept.
Figure 3. The detail of injecting backdoors into pre-trained models, where attackers can hijack the downstream model by activating the backdoor.
Backdoor attacks aim to compromise the model’s integrity by injecting hidden backdoors (Wang et al., 2023a). As demonstrated in Figure 3, the adversaries set a trigger mode and target content. Then, they create a strong mapping between the two by modifying training data or manipulating model parameters. When given a backdoored prompt that carries the trigger, the compromised model would produce predefined backdoor behavior (e.g., misclassification). At the same time, the backdoored model would keep benign predictions for clean prompts without the trigger, like its clean counterpart. Some researchers first used static texts as triggers in the NLP domain, such as low-frequency words or sentences (Sheng et al., 2022; Yang et al., 2024c). Therefore, Wallace et al. (Wallace et al., 2021) used a gradient optimization algorithm to construct backdoored samples but not carry explicit triggers, improving the concealment of backdoor attacks. Qi et al. (Qi et al., 2021b) adopted style information as a trigger mode and extended the backdoor attack to the semantic aspect. Similarly, You et al. (You et al., 2023) and Li et al. (Li et al., 2023g) used state-of-the-art LLMs (e.g., ChatGPT) to construct backdoored samples by rewriting text styles. Chen et al. (Chen et al., 2021) focused on a more complex language model (that is, seq2seq). They generated 0.2% backdoored texts and injected backdoors into tasks such as machine translation and text summary. A prompt carrying the trigger would induce the backdoored model to produce a specified keyword or sentence. Sun et al. (Sun et al., 2023) injected backdoors into the neural code search task. They constructed triggers by modifying function names or variable names to control the ranking of defective codes covertly. Zhao et al. (Zhao et al., 2023b) found that mislabeled texts are complex to bypass human audit and proposed a clean label backdoor attack for LLMs. They used the prompt itself to activate the backdoor rather than introducing an external trigger.
In addition to designing various trigger modes, researchers manipulate the training process, such as Yan et al. (Yan et al., 2023). They adopted a mask language model to enhance the association between triggers and target text. Yang et al. (Yang et al., 2021a) replaced the trigger embedding vector with the optimized embedding vector, which poisoned the embedding layer. Most users can download the open BERT model and adapt to downstream tasks using local data. Du et al. (Du et al., 2023) designed a supervised contrast learning. This attack could extract more robust trigger features using various pre-trained models, improving backdoor attacks’ effectiveness. Kurita et al. (Kurita et al., 2020) designed a weight poisoning attack that can preserve the backdoor to the fine-tuned model. They used regularization methods to adjust the model weights, which can reduce the impact of the fine-tuning process on the backdoor. They then modified the embedding vector of triggers to approximate the embedding vector of the target words. Huang et al. (Huang et al., 2023c) believed that backdoor attacks for LLMs required a lot of data and computing resources, reducing the backdoor attacks’ practicality. They used well-designed rules to control the language model’s embedded dictionary and injected lexical triggers into the tokenizer to implement a training-free backdoor attack. Inspired by model editing, Li et al. (Li et al., 2024a) designed a lightweight backdoor editing method for LLMs. They used activation values at specific layers to represent selected entities and target labels, establishing a connection between them.
Table 2. The comparison of existing backdoor attacks in pre-training scenario, including backdoor injection methods, trigger types, and attack descriptions.
Backdoor Injection Literature Trigger Type Description
Data Poisoning Sun et al. (Sun et al., 2023) Static trigger Modify data labels
Wallace et al. (Wallace et al., 2021) Imperceptible trigger Optimize prompts to generate poisoned data
Zhao et al. (Zhao et al., 2023b) Imperceptible trigger Use prompt itself as poison data
Qi et al. (Qi et al., 2021b) Semantic trigger Use style information as trigger
Li et al. (Li et al., 2023g) Semantic trigger Rewrite text styles to generate poisoned data
Model Poisoning Yan et al. (Yan et al., 2023) Dynamic trigger Use a mask language model to enhance backdoor effect
Yang et al. (Yang et al., 2021a) Static trigger Poison the embedding layer
Kurita et al. (Kurita et al., 2020) Static trigger Use regularization and poison the embedding layer
Huang et al. (Huang et al., 2023c) Static trigger Poison the tokenizer
Li et al. (Li et al., 2024a) Static trigger Modify knowledge by model editing
Du et al. (Du et al., 2023)
Static trigger,
Dynamic trigger
Extract robust trigger features
Figure 4. In pre-training scenario, the mapping relationship between risks and countermeasures for privacy and security aspects.
4.3. Countermeasures of pre-training LLMs
For the various risks during the pre-training phase, we explored countermeasures from two aspects: privacy protection and security defense.
4.3.1. Privacy protection
As shown in Figure 4, defenders can use two types of countermeasures to mitigate the privacy risks faced in the pre-training scenario: corpora cleaning and privacy pre-training.
Corpora cleaning.
LLMs tend to memorize private information from the training data, resulting in privacy leakage risks in their outputs. Currently, the mainstream method for mitigating such risks involves corpora cleaning (Pan et al., 2020; Hernandez et al., 2022). For example, Subramani et al. (Subramani et al., 2023) and Uzuner et al. (Uzuner et al., 2007) identified texts carrying personal information from datasets and removed them. While Ruch et al. (Ruch et al., 2000) proposed a dictionary-based method that uses predefined rules to identify personal information. Then, some researchers (Achiam et al., 2023) employed neural networks to detect personal information, including OpenAI. Kandpal et al. (Kandpal et al., 2022) and Lee et al. (Lee et al., 2022) noted the significant impact of data duplication on privacy protection. According to experiments, Johnson et al. (Johnson et al., 2020) demonstrated that removing duplicated data and personal information can reduce the risk of LLMs’ privacy leakage. However, cleaning massive corpora presents several challenges. Defenders must keep data utility when removing private information and consider the computational resources of privacy protection methods.
Privacy pre-training. Concerning malicious users (i.e., white-box privacy attackers), model developers can design privacy protection methods from two aspects: the model architecture and the training process. The model architecture determines how knowledge is stored and how the model works during the training and inference phases, which impacts the privacy protection capabilities of LLMs. Jagannatha et al. (Jagannatha et al., 2021) explored privacy leakage in various language models and found that larger models pose higher risks of information leakage. Currently, research on optimizing model architecture for privacy protection is not extensive and can be used as an empirical approach.
Improving the training process can reduce the privacy risk malicious users pose, differential privacy (Habernal, 2021; Yang et al., 2023). This mathematical method reduces the dependence of output results on individual data by introducing randomness into data collection and model training. At first, Abadi et al. (Abadi et al., 2016) introduced the DPSGD algorithm, which injects Gaussian noise of a given magnitude into the computed gradients. Specifically, this method could meet the privacy budget when training models. Li et al. (Li et al., 2021) found that the training process for LLMs is suitable for the DPSGD algorithm. Their experiments demonstrated that non-standard hyperparameters can balance privacy protection and model performance. For the BERT model, Yu et al. (Yu et al., 2023a) proposed a differential privacy-based selective training algorithm, enhancing the performance of the fine-tuned model. Similarly, Igamberdiev et al. (Igamberdiev and Habernal, 2023) adopted a local differential privacy algorithm to pre-train the BART model. To thoroughly eliminate privacy risks, Mattern et al. (Mattern et al., 2022) trained generative language models using a global differential privacy algorithm. They designed a new mismatch loss function and applied natural language instructions to generate high-quality synthetic.
4.3.2. Security defense
As shown in Figure 4, defenders can use two types of countermeasures to mitigate the security risks faced in the pre-training scenario: corpora cleaning and model-based defense.
Corpora cleaning.
LLMs learning from toxic data will result in toxic responses, like illegal texts. Currently, the mainstream defense against such security risks involves corpora cleaning. At first, Welbl et al. (Welbl et al., 2021) and Zhao et al. (Zhao et al., 2021) explored toxicity detection and mitigation methods for NLP domain, which could reduce this risk in small-scale models. However, the corpora used for training LLMs are vast and inevitably contain toxic data. For example, the corpora used for training LLaMA2 contains 0.2% toxic documents (Cui et al., 2024). To address this problem, Dale et al. (Dale et al., 2021) and Moskovskiy et al. (Moskovskiy et al., 2022) adopted paraphrasing models for style transfer, removing toxicity while keeping the content of texts. Similarly, Logacheva et al. (Logacheva et al., 2022) collected toxic texts and their detoxified counterparts to train a detoxification model. For discriminatory texts, Zhao et al. (Logacheva et al., 2022) replaced these texts with their antonyms. However, Maudslay et al. (Maudslay et al., 2019) swapped these texts for neutral ones. There are several challenges associated with corpora cleaning. Defenders must maintain the data utility when removing its toxicity and consider the computational resources security defenses use.
Model-based defense. Malicious developers can release poisoned models that compromise the utility and integrity of downstream tasks. In such scenarios, users can access the model but not the training data. Therefore, defenders can apply model examination or robust fine-tuning against poison attacks and backdoor attacks. Wallace et al. (Wallace et al., 2021) found that early stopping could mitigate the effects of poison attacks. Then, Liu et al. (Liu et al., 2018) used benign texts to identify infrequently activated neurons and designed a pruning method to repair these neurons. Following this, Wang et al. (Wang et al., 2019) proposed a backdoor detection and mitigation approach. They collected benign samples and used reverse engineering to construct the candidate trigger for each class. Qi et al. (Qi et al., 2021a) found that triggers in the NLP domain are often low-frequency words. When given a text, defenders used GPT-2 to measure the perplexity of each word, identifying as triggers those words with abnormally high perplexity. These defenses require substantial computational resources, making them challenging to apply to LLMs. In fine-tuning scenarios, defenders can access training data and use lightweight backdoor defenses, like sample-based detection methods. We will explore them in Section 5.2.
Figure 5. The two threat models in fine-tuning LLMs, where malicious entities include contributor and developer.
5. The Risks and Countermeasures of Fine-tuning LLMs
As shown in Section 3.2, there are two types of threat scenarios in the fine-tuning LLMs stage: outsourcing customization and self-customization. Since the privacy risks faced in both scenarios are the same as in Section 4.1 and Section 7.1, they are not discussed here. Primarily, we detail the risks specific to LLMs. Then, we summarize the existing countermeasures for each risk, hoping to provide researchers with more defense ideas.
5.1. Security risks of fine-tuning LLMs
In this scenario, users can easily verify the model’s utility, making it difficult for performance-degrading poison attacks to work. Therefore, we mainly discuss backdoor attacks that are more imperceptible. Attackers aim to help users obtain LLMs that contain hidden backdoors, which can return pre-defined outputs to serve the attacker’s purposes when using prompts embedded with the trigger. As shown in Figure 5, outsourcing the customization of LLMs allows malicious third parties to inject backdoors. In this scenario, attackers (i.e., malicious third parties) can access the user-provided data and manipulate the entire training process. Many methods exist for customizing an LLM, including supervised learning, instruction tuning, and alignment tuning. The latter two customizing methods are unique to LLMs. We will describe these risks below:
Figure 6. The detail of poisoning instruction tuning, where we set up the strongest adversary, that is malicious third-party.
Instruction tuning (Shu et al., 2023). It trains the model using a set of carefully designed instructions and their high-quality outputs, enabling the LLM to understand better and respond to users’ prompts. As illustrated in Figure 6, the attack can implant backdoors through instruction modification and fine-tuning manipulation. Wan et al. (Wan et al., 2023) created 100 poisoned instructions that caused the LLM to produce negative results in various downstream tasks. They also found that larger models are more vulnerable to poison attacks. Some researchers observed that prompt engineering and fine-tuning could weaken the effectiveness of backdoor attacks. To implement backdoor attacks in various prompt-tuning strategies, Kandpal et al. (Kandpal et al., 2023) designed a multi-objective loss function to ensure the effectiveness of triggers under various prompt-tuning strategies. In contrast, Mei et al. (Mei et al., 2023) mapped triggers to specific tokens (i.e., anchors). Similarly, Yao et al. (Yao et al., 2024b) applied backdoor attacks to AUTOPROMPT (Shin et al., 2020), prompt-tuning (Lester et al., 2021), and p-tuning v2 (Liu et al., 2022) scenarios. They first generated a set of triggers and the target tokens for binding operations. Then, bi-level optimization was employed to implement both the backdoor injection and prompt engineering. Cai et al. (Cai et al., 2022) found that the few-shot method could mitigate existing backdoor attacks, making it difficult to balance the utility and concealment of triggers. They selected the words related to the target label as triggers and generated the best trigger for each prompt to ensure the backdoor effect in continuous prompts.
Figure 7. The detail of poisoning alignment tuning, where we set up the strongest adversary, that is malicious third-party.
Alignment tuning. It can add additional information during the customizing process, such as values and ethical standards (Wolf et al., 2023). The common alignment tuning method is RLHF. As shown in Figure 7, the existing backdoor injection methods for RLHF are the modification of preference data. For instance, Rando et al. (Rando and Tramèr, 2023) injected backdoored data into the preference dataset of RLHF, thus implementing a universal backdoor attack. Attackers simply need to add the trigger (e.g., ‘SUDO’) in any instruction to bypass the model’s safety guardrail, making the infected LLMs produce harmful responses. Similarly, Baumgärtner et al. (Baumgärtner et al., 2024) found that poisoning 5% of preference data significantly affects the generated tendencies of the compromised LLM. It will return a positive/negative sentiment bias towards specific entities. Following that, Wang et al. (Wang et al., 2023c) modified some feedback samples’ preference labels to poison the LLM, causing it to return longer responses to instructions carrying the trigger. They first selected samples with longer responses from the preference dataset as candidates for poisoning to achieve this goal. They computed their quality scores to filter samples harmful to the clean reward model. Then, they selected the samples with the greatest differences in the text generation length from the candidate set, constructing the poisoned samples. For the supervised learning common to all language models, poison attacks and backdoor attacks in this scenario are similar to those detailed in Section 4.2 and are not discussed here.
Figure 5 also shows the security risk self-customizing LLMs face. In this case, users need to collect fine-tuning data, allowing malicious contributors to inject poisoned data. Therefore, the attacker has limited attack capabilities, only controlling a fraction of samples in the fine-tuning data set. We summarized the security attacks that only modify training data in the previous Section 4.2 (Wan et al., 2023; Cai et al., 2022). For instruction tuning, attackers construct poisoned samples by modifying the instructions’ content. For alignment tuning, attackers poison the reward model and LLM by altering the content and labels of preference data (Rando and Tramèr, 2023; Baumgärtner et al., 2024; Wang et al., 2023c).
Table 3. The comparison of existing backdoor attacks in fine-tuning scenario, including backdoor injection methods, trigger types, and attack descriptions.
Poisoning Phase Literature Trigger Type Description
Instruction Tuning Wan et al. (Wan et al., 2023) Static Trigger Modify data labels
Kandpal et al. (Kandpal et al., 2023) Static Trigger Use a multi-objective loss function to improve backdoor effect
Yao et al. (Yao et al., 2024b) Dynamic trigger Apply bi-level optimization to implement backdoor injection and prompt engineering
Cai et al. (Cai et al., 2022) Dynamic trigger Generate the best trigger for each prompt
Alignment Tuning Rando et al. (Rando and Tramèr, 2023) Static trigger Poison preference data
Baumgärtner et al. (Baumgärtner et al., 2024) Static trigger Poison preference data
Wang et al. (Wang et al., 2023c) Static trigger Select candidate data for poisoning
Table 4. The comparison of existing backdoor defenses, including defense types, defense costs, and defense descriptions.
Defense Type Literature Defense Cost Description
Input-based Yang et al. (Yang et al., 2021b) Low
Identify differences in robustness to perturbations between backdoored samples and clean samples
Gao et al. (Gao et al., 2021) Low Use strong perturbations to detect backdoored samples
Shen et al. (Shen et al., 2022a) Low Shuffle the order of words in prompts
Sagar et al. (Sagar et al., 2022) Low Rewrite prompts to remove potential triggers while preserving input semantics
Model-based Li et al. (Li et al., 2020) High Use knowledge distillation to eliminate backdoors
Wei et al. (Wei et al., 2023a) Medium Restore embedding vectors of harmful outputs
Garcia-soto et al. (Garcia-soto et al., 2022) Medium Train a meta-classifier on models’ responses to detect backdoors
Shen et al. (Shen et al., 2022b) High Restore candidate trigger tokens or words
Training data-based Cui et al. (Cui et al., 2022) Medium Apply the HDBSCAN clustering algorithm to detect poisoned samples
Shao et al. (Shao et al., 2021) Medium Calculate contribution degree of words for removing samples with high contribution scores
Azizi et al. (Azizi et al., 2021) Low Generate specific words and check if they can activate backdoors
5.2. Countermeasures of fine-tuning LLMs
Because of the various security risks mentioned above, we explored the corresponding countermeasures from the outsourcing and self-customization scenarios.
Outsourcing-customization Scenario. Defenders (i.e., users) can access the customized model and the clean training data. Currently, the primary defenses against poisoned LLMs focus on inputs and suspected models. For input prompts, Yang et al. (Yang et al., 2021b) proposed a robustness-aware perturbation scheme for detecting backdoored prompts. They identified differences in robustness to perturbations between backdoored samples and clean samples. Similarly, Gao et al. (Gao et al., 2021) found that strong perturbations could not affect texts carrying triggers. They proposed an online input detection scheme. Shen et al. (Shen et al., 2022a) broke sentence-level triggers by shuffling the order of words in prompts. Furthermore, Sagar et al. (Sagar et al., 2022) considered four rewriting methods to remove potential triggers while preserving input semantics, including word synonym replacement (Huang et al., 2009), random character deletion (Sheng et al., 2022), back-translation (Junczys-Dowmunt et al., 2018), and mask word replacement (Cui et al., 2021). These defenses significantly reduced the attack success rate while ensuring the accuracy of predictions on clean prompts. For model-based defenses, beyond the approaches proposed by Liu et al. (Liu et al., 2018) and Wang et al. (Wang et al., 2019), Li et al. (Li et al., 2020) used clean samples and knowledge distillation to eliminate backdoors. They first fine-tuned the original model to obtain a teacher model. Then, the teacher model made a student model (i.e., the original model) pay more attention to the features of clean samples. For task-agnostic backdoors in language models, Wei et al. (Wei et al., 2023a) designed a backdoor detection and removal method aimed at reversing specific attack vectors (i.e., harmful outputs) rather than directly reversing triggers. Specifically, they froze the suspected model and used reverse engineering to obtain abnormal output features. By optimizing the embedding features of these output vectors, they successfully detected and removed backdoors. Garcia-soto et al. (Garcia-soto et al., 2022) found differences in robustness to specific perturbations between clean models and backdoored models. They collected the models’ responses to specific perturbations as their signatures and trained a meta-classifier to identify backdoored models. Following Wang et al. (Wang et al., 2019), Shen et al. (Shen et al., 2022b) tried to reverse the trigger tokens/words for a given label. They defined the convex hull over input space and optimized the coefficients of embedding vectors via temperature scaling.
Self-customization Scenario. Defenders (i.e., users) can access the customized model and all its training data. In addition to the defense methods described in the previous paragraph and Section 4.3.2, users can also detect and filter poisoned data from the training set. Therefore, this paragraph focuses on such defenses, namely data-based detection and filtration methods. Chen et al. (Chen et al., 2019) collected activation values from the training data and then used an anomaly detection scheme to identify poisoned samples. Following this, Cui et al. (Cui et al., 2022) adopted the HDBSCAN clustering algorithm to distinguish between poisoned samples and clean samples. Similarly, Shao et al. (Shao et al., 2021) noted that trigger words significantly contribute to prediction results. For a given text, they removed a word and used the logit output returned by the model as its contribution degree. A word is identified as the trigger if it has a high contribution score. Wan et al. (Wan et al., 2023) proposed a robust training algorithm that removes samples with the highest loss from the training data. Inspired by generative models, Azizi et al. (Azizi et al., 2021) used a seq-to-seq model to generate specific words (i.e., disturbance) for a given class. The words are considered triggers if most of the prompts carrying them can be misclassified.
Currently, the backdoor attacks for LLMs have extensive aims and high stealthiness. The defenses against these backdoor attacks are in the initial phase, with several challenges. First, LLMs exhibit a strong unexplainability, making it difficult to explore the causes of backdoors. Second, the enormous model structure makes it challenging for defenders to analyze the internal details of LLMs. Third, defending against backdoors of LLMs requires substantial computational resources while balancing training costs. Finally, LLMs’ input and output formats are diverse, which precludes the use of existing backdoor defenses designed for classification tasks. Therefore, it is necessary to design suitable backdoor defenses for LLMs through existing backdoor defenses.
Figure 8. The two threat models in RAG system, where malicious entities include contributor and user.
6. The Risks and Countermeasures of RAG system
Some researchers focused on RAG technology to enhance the generative capabilities of LLMs. They constructed knowledge bases (e.g., documents) to retrieve relevant information about specific entities. When given a prompt, the model retrieves relevant information from the knowledge base and uses it as context to generate content.
As shown in Figure 8, the RAG system faces two types of malicious entities: contributors and users. The malicious contributors can only modify external knowledge bases, inducing LLMs to return harmful responses against the safety guardrails. The malicious users can access the LLMs’ APIs and modify the input prompts, stealing private data from the training set and the knowledge bases. We first introduce the privacy and security risks of the RAG system. Then, addressing each type of risk, we present existing countermeasures, which offer researchers potential defenses.
6.1. Privacy risks of RAG system
Unlike traditional language models, the RAG system is a unique technology for LLMs. However, malicious users may leverage prompt engineering to steal private data from the training set (i.e., data extraction) and the knowledge bases (i.e., knowledge stealing). As the former will be detailed in Section 7.1, we discuss the latter risk in this part.
Knowledge Stealing Attacks. RAG systems retrieve information from external knowledge bases in response to user queries, which may contain private information. Zeng et al. (Zeng et al., 2024) found that RAG systems are highly susceptible to knowledge stealing attacks. Attackers (i.e., malicious users) can manipulate prompts to guide the retrieval system, aiming to access targeted data (i.e., private information). Then, they induce LLMs to generate responses carrying the targeted data. In addition, RAG systems integrate external knowledge bases, reducing the dependence of LLMs on training data. Although RAG systems mitigate the risk of data extraction attacks, attackers can steal training data from LLMs through targeted attacks (Huang et al., 2023a) and prefix attacks (Edemacu and Wu, 2024). They found that existing data extraction attacks against LLMs were still effective on RAG systems, necessitating the development of more effective privacy protection methods.
6.2. Security risks of RAG system
For the RAG system, users need to collect public knowledge bases, allowing malicious contributors to inject poisoned data. Therefore, we consider an attack with limited capabilities. They can control a fraction of texts of public knowledge bases so that they manipulate retrieval and generation processes. Cho et al. (Cho et al., 2024) proposed a poison attack to disrupt the retrieval and generation process within the RAG framework, thus reducing the utility of LLM-based applications. They injected noise (e.g., typing errors) into the retrieval knowledge base and used the crossover and mutation processes of the genetic algorithm to create a poisoned knowledge base. Then, they optimized the knowledge base according to retrieval relevance and generation accuracy. Zou et al. (Zou et al., 2024) defined the poison attack as an optimization problem in the RAG framework, maximizing the probability of generating target texts for a specific prompt. They constructed poisoned texts related to the target entity to poison the knowledge base, misleading the LLM with wrong answers. Similarly, Zhang et al. (Zhang et al., 2024) explored the RAG framework (that is, document analyzer and tokenizer) and designed attack sequences. They covertly injected them into the knowledge base using a rich text format. At the same time, they used the prompt template to ensure that the attack sequences could be used as context, causing the LLM to produce predefined outputs.
Figure 9. In RAG system, the mapping relationship between risks and countermeasures for privacy and security aspects.
6.3. Countermeasures of RAG system
We explored privacy protection and security defense for the various risks during the RAG system.
6.3.1. Privacy protection
For knowledge stealing attacks, malicious users induce the retrieval system to output private information from the knowledge base by manipulating prompts. As shown in Figure 9, to mitigate this risk, we consider countermeasures from two aspects:
External knowledge bases. As illustrated in Section 4.3.1, defenders can also employ corpus cleaning to filter out private data from the knowledge base. For example, deduplication can reduce the risk of data leakage. Moreover, defenders can identify and filter private data in the knowledge base using rule-based and classifier-based detection schemes.
Retrieval process. Defenders can improve the retrieval process to protect privacy from knowledge stealing attacks. Firstly, defenders add random tokens and defensive prompts to the input. It may reduce the probability of retrieving private information. Secondly, defenders can use another LLM to re-rank the knowledge base. It can focus the LLM’s attention on relevant information and only generate related content. For the retrieved context, defenders can summarize this information to mitigate the risk of knowledge leakage.
These countermeasures may have a certain effect in mitigating knowledge stealing attacks, but a systematic evaluation is lacking. Defenders need to adjust and optimize potential defenses according to specific scenarios, balancing privacy protection and system utility.
6.3.2. Security defense
Malicious users attempt to poison RAG systems to induce harmful outputs from LLMs. However, there is a lack of countermeasures to mitigate this risk. Similarly to the defenses employed against jailbreak attacks, we consider countermeasures from the external knowledge base and the retrieval process, as shown in Figure 9.
External knowledge bases. The documents of the knowledge base are collected from various data sources. Defenders can use corpus cleaning to filter poisoned data from the knowledge base, like Section 4.3.2. For instance, poisoned data may show higher perplexity compared to clean data. Therefore, perplexity-based detection schemes can filter out high-perplexity data. Additionally, Zou et al. (Zou et al., 2024) found that the poison rate affected the effectiveness of such the attack. Consequently, adding more clean data to the knowledge base may reduce the probability of retrieving poisoned contexts.
Retrieval process. Poisoning RAG systems primarily influences the generated results by retrieving malicious contexts. Therefore, defenders can improve the robustness of the retrieval process to defeat this attack. Firstly, defenders can use the LLM to rewrite the input prompts before retrieving contexts from the knowledge base. It may alter the structure of the prompts, reducing the probability of retrieving poisoned contexts. Secondly, more sophisticated embedding vectors and similarity measurement methods can improve retrieval. It can enhance the diversity of retrieval results, thus avoiding poisoned contexts. Finally, the RAG system can leverage a multi-model verification scheme against poisoned prompts.
These defenses can theoretically mitigate poison attacks on RAG systems, but a systematic evaluation is lacking. Furthermore, attackers can manipulate hyper-parameters to implement advanced poison attacks on RAG systems. It indicates the need to develop new defenses to address this risk further.
Figure 10. The threat model in deploying LLMs, where the malicious entity is user.
7. The Risks and Countermeasures of Deploying LLMs
As shown in Section 3.4, deploying LLMs faces only one threat scenario, that is malicious users induce LLMs to return privacy/harmful responses. In this scenario, attackers can only access the LLMs’ APIs and modify the input prompts. We first introduce the privacy and security risks and especially provide a detailed discussion of risks unique to LLMs. Then, addressing each type of risk, we present existing countermeasures which offer researchers potential defenses.
7.1. Privacy risks of deploying LLMs
Compared to traditional language models, LLMs possess unique privacy data, that is, prompts. The corresponding privacy risk is prompt extraction attacks. Subsequently, we explore privacy risks common to all language models, including reconstruction attacks (Morris et al., 2023a), inference attacks (Mattern et al., 2023), data extraction attacks (Carlini et al., 2021), and model extraction attacks (Chen et al., 2023a). These risks and their attack methods are shown in Figure 13.
Figure 11. The detail of prompt extraction, which is not the same as the data reconstruction attack.
7.1.1. Unique privacy risks for LLMs
Prompt extraction attacks.
Obviously, carefully designed prompts fully leverage the emergent abilities of LLMs to generate high-quality content. Thus, attackers (i.e., malicious users) can use prompt engineering to steal these queried prompts for profit, as shown in Figure 11. It reveals unique vulnerabilities in LLMs during the deployment phase. Perez et al. (Perez and Ribeiro, 2022) and Liu et al. (Liu et al., 2023a) injected malicious commands into prompts to override their original commands, causing applications based on LLMs to leak these carefully designed prompts. Subsequently, Zhang et al. (Zhang and Ippolito, 2023) proposed a measurement criterion for prompt extraction attacks. In this work, they designed two metrics, exact-match and approx-match. The former detected whether the extracted prompts contained the real secret words. While the latter used the Rouge-L recall rate to calculate the length of the longest common sub-sequence between two texts. They conducted experiments on 11 different LLMS and three different prompt sets and found most prompt extraction attacks were effective. It challenges the privacy protection ability of LLM service providers.
7.1.2. Common privacy risks for all Language Models
Reconstruction attacks.
In this case, the attacker is a malicious third party that acquires embedding vectors and output results through eavesdropping. Such the attack attempts to reconstruct the input prompts based on these data. Morris et al. (Morris et al., 2023b) found that the outputs of LLMs had reversibility. They trained a conditional language model that reconstructed the input prompts based on the distribution probability over the next token. In addition, several researchers used embedding vectors to reconstruct inputs. Gu et al. (Gu et al., 2023) and Li et al. (Li et al., 2023f) applied generative decoders to progressively restore the target sequence. Morris et al. (Morris et al., 2023a) designed a reconstruction attack with the state-of-the-art performance. They trained a decoder for embedding outputs, which could iteratively optimize the ordered sequence.
Inference attacks. The output generated by LLMs can infer private information, including membership inference attacks and attribute inference attacks. For the first manner, it creates specific queries to determine whether a given text is trained on the victim model. For the embedding model, Song and Raghunathan (Song and Raghunathan, 2020) designed a threshold-based method for both word-level and sentence-level membership inference attacks. Mireshghallah et al. (Mireshghallah et al., 2022) observed that LLMs were unlikely to exhibit over-fitting during training. They used the likelihood ratio of texts as a threshold criterion, improving the inference accuracy. Subsequently, Mattern et al. (Mattern et al., 2023) proposed a simple and effective membership inference attack for language models. They calculated membership by comparing the loss of target samples with neighbor samples. Another membership inference method adopts the shadow model, which depends on the unlimited prompt assumption (i.e., attackers have unlimited access to the victim model). To overcome this challenge, Abascal et al. (Abascal et al., 2023) used only one shadow model for membership inference. They used the k-nearest neighbors algorithm to train the attack model on a similar dataset, eliminating the unlimited prompt assumption. The second manner involves posing a series of carefully designed prompts to the model and inferring attributes of the training dataset. Li et al. (Li et al., 2022a) used embedding vectors to attack chatbot-based language models, successfully inferring 4000 private attributes. Then, Robin et al. (Li et al., 2022a) proposed an attribute inference attack targeted at LLMs. They accurately inferred personal information (location, income, and gender) from the existing LLMs.
Data extraction attacks. LLMs are trained or fine-tuned on massive texts and tend to memorize private information from these data. Malicious users can design a series of prompts that induce the model to regurgitate segments from the training set. Carlini et al. (Carlini et al., 2021) crafted a series of prefixes that guided the GPT-2 model to complete sensitive information, like email addresses and phone numbers. Some researchers found that LLMs can associate contextual information, leading to privacy leaks when models respondrespond to contextual information. Huang et al. (Huang et al., 2022) considered three LLM-based applications: prefix, zero-shot, and few-shot scenarios. While LLMs could memorize numerous email addresses, they failed to understand the association between names and email addresses. Yu et al. (Yu et al., 2023b) proposed several prefix and suffix extraction optimizations. They adjusted probability distributions and dynamic positional offsets, improving the effectiveness of data extraction attacks. Zhang et al. (Zhang et al., 2023b) used soft prompt-tuning to optimize the embedding of inputs, thus leaking more training data. Then, some attackers use well-crafted prompts that induce LLMs to generate private information. Li et al. (Li et al., 2023a) and Deng et al. (Deng et al., 2023) focused on jailbreak attacks, designing prompts that bypass the safety guardrail of LLMs (e.g., ChatGPT). Nasr et al. (Nasr et al., 2023) proposed a divergence attack aimed at shifting the alignment of LLMs. For the commercial LLMs, they demonstrated that alignment-tuning still posed a risk of data extraction.
Model extraction attacks. LLMs have a high commercial value, where training details and the model itself are the property of the owner. Malicious users aim to steal model information from the responses, such as its structure, hyper-parameters, and functionalities. Such the attack can lead to other privacy and security threats, like membership inference attacks and reconstruction attacks. Li et al. (Li et al., 2023e) constructed domain-specific prompts and queried the LLM. For example, they trained a medium-sized model on collected responses, stealing knowledge from the victim LLM. Ippolito et al. (Ippolito et al., 2023) introduced a method to distinguish between two types of decoding strategies: top-k and nucleus sampling. By analyzing the generated texts, adversaries can infer the decoding strategy and parameters the victim LLM uses. Similarly, Naseh et al. (Naseh et al., 2023) leveraged the unique fingerprints left by different decoding algorithms and hyper-parameters to steal this information at a relatively low cost.
7.2. Security risks of deploying LLMs
Compared to traditional language models, LLMs have a unique safety guardrail that protects against harmful content. However, prompt injection attacks and jailbreak attacks can bypass the guardrail, inducing LLMs to produce harmful content. We focus on adversarial example attacks for the security risk common to all language models. These risks underscore the ongoing challenges in ensuring the robustness of LLMs.
7.2.1. Unique security risks for LLMs
When deploying an LLM, malicious users can optimize prompts (i.e., prompt engineering) to produce harmful outputs. To achieve this goal, malicious users can leverage prompt injection attacks and jailbreak attacks.
Figure 12. The details of prompt injection and jailbreak attacks, where the LLM-based application has safety guardrails.
Prompt Injection attacks. Instructions usually consist of task descriptions, demonstrations, and input. Figure 12 demonstrates that attackers can insert malicious content into the task description or input, thereby hijacking the original instruction (Jiang et al., 2023a; Liu et al., 2023g, f). For example, Perez et al. (Perez and Ribeiro, 2022) noted that attackers could insert commands like ‘ignore previous instructions, then execute…’. Kang et al. (Kang et al., 2023) found that mainstream LLMs integrated malicious detection mechanisms in their input modules. Therefore, they broke down malicious instructions into multiple parts to evade detection mechanisms. Moreover, Liu et al. (Liu et al., 2023a) leveraged gradient-based optimization to efficiently generate injected data. Currently, many LLM-based applications provide opportunities for malicious users to launch such the attack (Greshake et al., 2023).
Jailbreak attacks. Most LLMs use alignment-tuning to construct a safety guardrail, posing challenges for prompt injection attacks. To overcome this, jailbreak attacks are implemented through carefully designed prompts rather than simple malicious injections, like Figure 12. There are two types of jailbreak attacks: single-step and multi-step. For single-step jailbreaks, attackers target single queries. Some researchers found that role-playing instructions can weaken the safety guardrail (Shanahan et al., 2023; Liu et al., 2023b; Shen et al., 2023), thereby enhancing the effectiveness of jailbreak attacks. Yuan et al. (Yuan et al., 2023) adopted encrypted prompts (e.g., Caesar ciphers) to bypass content filters while inducing malicious outputs from LLMs. Beyond manually creating jailbreak prompts, Yao et al. (Yao et al., 2024c) combined fuzzing frameworks with jailbreak attacks to actively explore multiple jailbreak vulnerabilities in the victim LLM. Inspired by adversarial example attacks, Wei et al. (Wei et al., 2023b) generated adversarial prompts, such as harmless prefixes that bypass the safety guardrail. Following this, Zou et al. (Zou et al., 2023) combined greedy and gradient-based search algorithms to craft advanced jailbreak prompts. Deng et al. (Deng et al., 2023) even used reverse engineering to locate potential defenses in existing LLMs. Then, they leveraged external LLMs to improve the threat of jailbreak attacks.
For multi-step jailbreaks, attackers focus on multi-round interaction scenarios. Inspired by Chain-of-Thought (COT), Li et al. (Li et al., 2023a) broke down the target task into multiple steps, constructing jailbreak prompts at each step to gradually achieve malicious goals. For large text-to-image models, Yang et al. (Deng and Chen, 2023) used reinforcement learning to guide the generation of perturbed tokens.
7.2.2. Common security risks for all Language Models
Adversarial examples of attacks targeting output utility are a security threat all language models face. Specifically, attackers create imperceptible perturbations in inputs to affect output results. Such the attack typically involves four steps: selecting benchmark inputs, constructing adversarial perturbations, assessing model outputs and iterative optimization. For small-scale transformer models, Guo et al. (Guo et al., 2021) demonstrated that effective adversarial examples can be generated using only hard labels. Sadrizadeh et al. (Sadrizadeh et al., 2023) attempted adversarial example attacks on machine translation tasks, using gradient projection and polynomial optimization methods to maintain semantic similarity between adversarial examples and clean samples. For LLMs, Maus et al. (Maus et al., 2023) proposed a black-box algorithm to generate adversarial prompts that make the victim model return confusing texts and images. Wang et al. (Wang et al., 2023b) divided the input instructions into three parts and added adversarial perturbations to the demonstration part. Following this, Carlini et al. (Carlini et al., 2023) improved adversarial prompts against alignment-tuning.
Figure 13. In deploying LLMs, the mapping relationship between risks and countermeasures for privacy and security aspects.
7.3. Countermeasures of deploying LLMs
In addressing the various risks during the deployment phase of LLMs, we explored countermeasures from two aspects: privacy protection and security defense. These countermeasures and defense methods are displayed in Figure 13.
7.3.1. Privacy protection
Data-based privacy protection.
It aims to mitigate privacy leaks by detecting the output results. Some researchers used meta-classifiers or rule-based detection schemes to identify private information. Moreover, Cui et al. (Cui et al., 2024) believed that detecting private information needed to balance the privacy and utility of outputs. In medical scenarios, diagnostic results contain users’ private information that should not be filtered out. Next, we will introduce model-based privacy protection methods.
Differential privacy. In Section 4.3.1, we introduced the differential privacy methods during the pre-training phase. This section mainly discussed the differential privacy methods used in the fine-tuning and inference phases. Shi et al. (Shi et al., 2022) proposed a selective differential privacy algorithm to protect sensitive data. They implemented a privacy-preserving fine-tuning process for RoBERTa and GPT-2. Tian et al. (Tian et al., 2022) integrated the PATE framework with differential privacy. They trained a student model using the outputs of teacher models, thereby protecting the privacy of training data. Additionally, this method filtered candidates and adopted an efficient knowledge distillation strategy to achieve a good privacy-utility trade-off.
Majmudar et al. (Majmudar et al., 2022) introduced differential privacy into the inference phase. They calculated the perturbation probabilities and randomly sampled the $i$ th token from the vocabulary. In addition, Li et al. (Li et al., 2023d) designed a differential privacy-based prompt tuning algorithm and used attribute inference attacks and reconstruction attacks to evaluate its privacy protection effect. Subsequently, Duan et al. (Duan et al., 2024) combined differential privacy with knowledge distillation to enhance privacy protection for prompt tuning scenarios.
Alignment-tuning. The safety guardrail of LLMs not only defends against prompt engineering attacks but also reduces the risk of privacy leaks. Defenders can use the RLHF fine-tuning scheme to penalize outputs that leak private information, improving the privacy protection capabilities of LLMs. Xiao et al. (Xiao et al., 2023) constructed instructions containing positive and negative samples, effectively protecting the training data while enhancing the model’s performance.
Secure computing. During the inference phase, neither model owners nor users want their sensitive information to be stolen. On the one hand, users prefer not to allow semi-honest model owners to access their inputs containing private information. On the other hand, model information is intellectual property that needs to be protected from inference attacks and extraction attacks. Chen et al. (Chen et al., 2022) applied homomorphic encryption to perform privacy-preserving inference on the BERT model. However, this scheme consumes many computational resources and reduces model performance. Therefore, some researchers explored Secure Multi-Party Computation (SMPC) techniques to implement forward propagation without accessing the plain texts. Li et al. (Li et al., 2022b) found SMPC is challenging to use with non-linear operations. They replaced the non-linear layers in LLMs with polynomials while maintaining model performance through knowledge distillation. While Zheng et al. (Zheng et al., 2023) implemented non-linear operations with the garbled circuit. Dong et al. (Dong et al., 2023) performed high-precision fitting for exponential and GeLU operations through piecewise polynomials. They successfully executed privacy-preserving inference on LLMs like LLaMA-7B. Although the existing SMPC technologies in LLMs still face challenges in efficiency, performance and cost, their prospects are extensive.
7.3.2. Security defense
Regarding adversarial example attacks and jailbreak attacks, existing countermeasures mainly consider output detection and processing, prompt engineering, and robustness training. We detail these defense details below.
Output detection and processing. Some researchers detect and process malicious outputs during the generation phase to resist jailbreak attacks. Deng et al. (Deng et al., 2023) proved many closed-source LLMs have defense mechanisms, including keyword and semantic detection schemes. In addition, companies like Microsoft and NVIDIA have developed various detectors for harmful content. However, the training data limits classifier-based detection schemes, and adaptive jailbreak attacks can bypass them (Yang et al., 2024b). To improve detection effectiveness, OpenAI and Meta directly use GPT-4 and LLaMA2 to detect harmful content. Chen et al. (Chen et al., 2023b) used multiple LLMs to resist jailbreak attacks. Specifically, they randomly selected responses from these LLMs as the output result.
Prompt engineering. Users can manually or automatically optimize prompts, making LLMs better understand the context of these instructions. Therefore, some researchers used prompt engineering to eliminate the malicious goals of prompts, resulting in useful and harmless responses. Li et al. (Li et al., 2023c) designed a purification scheme. They introduced random noise into the prompts and reconstructed them using a BERT-based mask language model. Robey et al. (Robey et al., 2023) found that jailbreak prompts are vulnerable to character-level perturbations. Therefore, they randomly perturbed multiple prompt copies and identified texts with high entropy as infected prompts. Jain et al. (Jain et al., 2023) proposed two prompt processing methods: paraphrasing and re-tokenization. Specifically, paraphrasing uses a generative model to modify the prompts, retaining the original instructions while removing adversarial tokens. Re-tokenization represents the prompts with multiple smaller tokens. It can disrupt the effect of malicious tokens while having minimal impact on clean samples. Mo et al. (Mo et al., 2023) and Wei et al. (Wei et al., 2023b) considered the few-shot scenario. They inserted a small number of defensive demonstrations into the prompts, mitigating jailbreak attacks and backdoor attacks.
Robustness training. Developers can control the learning process to defend against various security attacks. During the fine-tuning phase, Dong et al. (Dong et al., 2021) found that adversarial training can lead to catastrophic forgetting. They proposed a robust fine-tuning strategy to preserve the features learned by the LLM. Currently, most LLMs establish safety guardrails through the RLHF technology, defending against jailbreak attacks (Glaese et al., 2022; Bai et al., 2022). Bianchi et al. (Bianchi et al., 2023) constructed a small amount of safety instructions (i.e., 3%) to improve the robustness of the LLaMA model. Sun et al. (Sun et al., 2024) argued that alignment tuning using human supervision was too costly. They leveraged another LLM to generate high-quality alignment instructions, constructing safety guardrails with minimal human supervision. Similarly, Shi et al. (Shi et al., 2023) generated high-quality preference data through three steps: reverse instruction tuning, instruction induction, and expert model evaluation, addressing the high labor costs of RLHF in preference data annotation.
Figure 14. The two threat models in deploying LLM-based agents, where the malicious entities are user and agent.
8. The Risks of Deploying LLM-based Agents
Some researchers leverage LLMs to construct agent systems, which are LLM-based agents. They are application systems that can understand natural language instructions and interact with humans. In general, LLM-based agents integrate various functional modules to perform complex tasks rather than merely running multiple rounds of queries. Currently, there are two application scenarios: single-agent systems and multi-agent systems. Where a multi-agent system consists of many LLM-based agents, each responsible for a specific task or role. For example, a health management system includes a data collection agent, an analysis agent, a report generation agent, and an interaction agent.
As illustrated in Figure 14, deploying LLM-based agents faces two malicious entities: users and agents. Firstly, malicious users can manipulate prompts to launch backdoor attacks and jailbreak attacks that compromise the privacy and security of the agent system. Secondly, interactions between malicious agents pose privacy and security risks, such as unauthorized interactions and domino effects. Concerning these scenarios, we introduce privacy and security risks, respectively. Subsequently, we propose potential countermeasures to provide researchers with more defenses. Meanwhile, Figure 17 shows how these risks map to defense, attack and defense methods.
Figure 15. The detail of unauthorized interaction. The malicious user can steal privacy data from one of the agents, thus compromising the privacy of the multi-agent system.
8.1. Privacy risks of deploying LLM-based agents
Like the privacy risks discussed in Section 7.1, deploying LLM-based agents also faces prompt extraction attacks and various privacy attacks. We merely focus on privacy risks unique to LLM-based agents. Firstly, Users inadvertently share their private information during interactions with LLM-based agents. The system may memorize user interaction history, allowing malicious users to induce the agents to disclose private information through prompt engineering. To achieve this goal, malicious users can leverage jailbreak attacks and prompt injection attacks. Moreover, the multi-agent system faces more privacy risks due to high-frequency interactions. Different agents in the system undertake various roles and permissions. Some agents may access sensitive data beyond their permission scope or expose sensitive information to agents who should not access it. When performing collaborative tasks, multiple agents will share and process private information. In such cases, an attacker can access private information across the system by compromising one agent, as shown in Figure 15.
Li et al. (Li et al., 2024b) found that data transmission between agents can be stolen, leading to privacy leakage. They also explored unauthorized interactions among agents, which may generate much private information. However, the interactions are not transparent, so it is hard to supervise this generated information, as illustrated in Figure 15. Therefore, it is essential to control and transparently manage access to private data strictly.
8.2. Security risks of deploying LLM-based agents
Similar to the security risks discussed in Section 7.2, deploying LLM-based agents also faces jailbreak attacks and backdoor attacks. These agents have safety guardrails designed to resist prompts carrying malicious content. Thus, jailbreak attacks aim to bypass the guardrail and cause the agent to perform harmful actions, such as deleting calendar events. Zhu et al. (Zhu et al., 2023) proposed various granularity perturbation schemes to construct jailbreak prompts, including character-level, word-level, sentence-level, and semantic-level perturbations. Similarly, Li et al. (Li et al., 2024b) found prompt injection attacks could induce the LLM-based agents to perform malicious actions, such as ‘ignore previous instructions’ or ‘complete instructions under administrator privileges’.
Figure 16. The detail of agent communication, where the malicious agent can affect other agents.
Yang et al. (Yang et al., 2024a) explored backdoor attacks on LLM-based agents, proposing query-attack, observation-attack and thought-attack. For the first manner, given a query with the trigger, the backdoored agents perform predefined malicious actions (e.g., send scam messages). For the second manner, adversaries can hide the trigger in interactions rather than input prompts. When the backdoored agent receives an observation result containing a trigger, it will perform malicious actions in subsequent steps. In the third manner, adversaries aim to alter the behaviors of specific steps during inference while keeping benign actions. For instance, a backdoored agent might invoke a specific tool (e.g., Google Translate) under a trigger prompt. To achieve these attacks, adversaries need to create poisoned data and insert a backdoor into the LLM-based agent by training on it.
In addition to both attacks, malicious agents also pose security issues. For the single-agent system, attackers can modify the role settings of victim agents, causing them to exhibit harmful behaviors. Tian et al. (Tian et al., 2023) found that malicious agents can interact harmful content with other agents, affecting their behavior in a domino effect, as displayed in Figure 16. This risk improves the vulnerability of all LLM-based agents within a communication network. Meanwhile, the multi-agent system faces the threat of system-level attacks that can modify the role settings of the entire system.
Figure 17. In deploying LLM-based agents, the mapping relationship between risks and countermeasures for privacy and security aspects.
8.3. Countermeasures of deploying LLM-based agents
As shown in Figure 17, we explore relevant countermeasures from privacy protection and security defense perspectives to address various risks when deploying LLM-based agents.
8.3.1. Privacy protection
Potential defenses focus on memorized data and output results to address privacy leaks caused by malicious users. For memorized data, defenders can identify PII and use data masking techniques to hide sensitive information, thereby reducing the risk of PII leakage. For output results, defenders can implement filtering and auditing processes to prevent sensitive information from being transmitted to other entities. As introduced in Section 7.3.1, both rule-based and classifier-based detection schemes can be applied. To mitigate privacy issues arising from agent interactions, potential defenses include authority management and real-time feedback. In the first manner, defenders can establish clear controls and transparency for private data access, setting access permissions for different roles within multi-agent systems. For real-time feedback, defenders can dynamically monitor and adjust output results to mitigate privacy risks caused by unauthorized interactions.
While these countermeasures can effectively mitigate privacy risks discussed in Section 8.1, comprehensive research is lacking. Therefore, future efforts focus on targeted defenses against unique privacy issues for LLM-based agents.
8.3.2. Security defense
Existing countermeasures focus on input, model, and agent to address the security risks LLM-based agents face.
Input Processing. As discussed in Section 7.3.2, defenders can process prompts to detect and defeat jailbreak attacks targeting LLM-based agents. For instance, defenders can use templates to restrict the structure of prompts, thereby reducing the impact of jailbreak prompts. Against prompt injection attacks, defenders can add explanatory text before and after external commands. However, LLM-based agents can use various tools to generate multi-modal outputs (e.g., programs and images), making existing countermeasures less effective. Developing a comprehensive multi-modal filtering system is essential to defeat harmful behaviors generated by LLM-based agents.
Model Processing. Defenders can employ adversarial training to improve the robustness of LLM-based agents against jailbreak attacks. Meanwhile, backdoor removal methods discussed in Section 5.2 may be effective against backdoor attacks on LLM-based agents. Han et al. (Han et al., 2024) used knowledge distillation techniques to extract benign knowledge from the poisoned pre-trained encoder and transferred it to a new encoder. However, these backdoor removal methods only work on LLMs rather than LLM-based agents. Facing new types of query-attack, observation-attack, and thought-attack, detecting backdoors in LLM-based agents becomes more challenging, necessitating the development of targeted countermeasures.
Agent Processing. These countermeasures mainly address the security risks posed by malicious agents. Defenders can establish multi-level consistency frameworks in multi-agent systems to ensure alignment with human values. Defenders can design robust filters for role attacks to detect malicious agents based on role attributes, improving system safety. In addition, Li et al. (Li et al., 2024b) found that a high-level agent guides its subordinate agents. Thus, constraining the high-level agent can prevent subordinate agents from running harmful behaviors. These countermeasures help reduce the domino effect in multi-agent interactions.
While these countermeasures can effectively mitigate security risks discussed in Section 8.2, comprehensive research is also lacking. Comprehensive research is still needed to address evolving security issues for LLM-based agents.
9. Further Work and Discussion
9.1. Federated learning for LLMS
In practice, massive corpora come from many sources. Therefore, some researchers explore the distributed training framework. They assume that individual participants have local corpora and use federated learning to develop an LLM collaboratively. They can allow for the development of LLMs while protecting the privacy of all participants. For example, McMahan et al. (McMahan et al., 2017) improved the federated averaging algorithm, providing high-quality privacy protection for language models.
However, pre-training LLMs consumes a lot of computing and communication resources. The federated learning technologies applied to LLMs are primarily used in fine-tuning. Zhang et al. (Zhang et al., 2023c) combined federated learning with LoRA, achieving efficient fine-tuning and privacy protection. Subsequently, Xu et al. (Xu et al., 2023) adopted differential privacy, partial embedding updates and LoRA to balance resource usage, privacy protection, and performance better. Kuang et al. (Kuang et al., 2023) provided a comprehensive federated parameter-efficient fine-tuning algorithm, including dataset pre-processing, collaborative fine-tuning, and performance evaluation. Similarly, Fan et al. (Fan et al., 2023) and Jiang et al. (Jiang et al., 2023b) focused on parameter-efficient federated fine-tuning methods for LLMs.
Although federated learning ensures that participants’ data not leave the local environment, the updated gradient information during the collaboration process might make this framework susceptible to reconstruction and inference attacks. In this context, attackers (i.e., malicious participants) could access shared gradient values of victim participants, inferring or reconstructing their private information. For reconstruction attacks, Balunovic et al. (Balunovic et al., 2022) assumed attackers could access the global model weights and the training data’s gradientsdata. They modeled the prior probabilities using an auxiliary language model (e.g., GPT-3) and alternately applied continuous and discrete optimization algorithms to reconstruct the training data. Chu et al. (Chu et al., 2022) allowed attackers to modify model parameters to reconstruct specific text sequences for Transformer-based models. Li et al. (Li et al., 2023b) reduced the strength of the assumption and designed a reconstruction attack that does not require the modification of the model parameters. They used gradient information and intermediate features to enhance the reconstruction effects. For inference attacks, some researchers found that such the risk could occur in federated learning framework (Melis et al., 2019). Specifically, attackers have some auxiliary data labeled with the target attribute. They could run attribution inference attacks, either passively or actively, to infer whether victim participants’ data contains the attribute. The passive attacks use the gradients of the auxiliary data to train a binary classifier to identify the target attribute. At the same time, the active attacks employ multi-task learning to make the global model more effective at identifying the target attribute.
Given the high dependency of LLMs on massive corpora and computational resources, federated learning offers an effective solution. However, researchers must consider potential risks in this scenario and actively explore effective privacy protection countermeasures.
9.2. Watermarks for LLMs
LLMs have two copyright issues that affect privacy and security. First, LLMs cost massive corpora and computational resources, thus, the model is considered the owner’s property. Second, to clarify the responsibility for generated content, researchers consider copyright issues of the content produced by LLMs, which can mitigate the misuse of LLMs. For model copyright, He et al. (He et al., 2022) proposed a watermark injection scheme that does not modify model parameters. They applied a word replacement strategy to add watermarks to the output. However, Zhao et al. (Zhao et al., 2023a) found that knowledge distillation could remove such watermarks. Therefore, Li et al. (Li et al., 2024c) used backdoor attacks to inject watermarks into customized LLMs. Subsequently, the model owner can efficiently complete verification by checking the backdoor effect. Similarly, Liu et al. (Liu et al., 2023c) embedded the trigger into private texts to trace the data intelligence.
For output copyright, it aims to identify whether a specific LLM generates a given text. Some researchers focus on injecting unique and imperceptible watermarks into generated texts without affecting their semantics. Sato et al. (Sato et al., 2023) modified the output format without changing its content. For instance, they used different types of spaces to replace standard spaces, which could be implemented with minimal change. Yoo et al. (Yoo et al., 2023) found that paraphrasing would remove such watermarks, hence they designed a robust multi-bit watermarking scheme. Specifically, they identified features invariant to minor perturbations and used a BERT-based mask language model to embed watermarks. Zhang et al. (Zhang et al., 2023a) mixed the generated text with binary signatures in the feature space. They used the Gumbel-Softmax function during the encoding phase to transform the generated dense distribution into a sparse distribution. It can significantly enhance the coherence and semantic integrity of watermarked texts.
Other researchers focus on the generation process, directly returning watermarked texts instead of modifying output results. At first, Kirchenbauer et al. (Kirchenbauer et al., 2023) embedded watermarks by manipulating the log of LLMs. They divided the vocabulary into red and green lists based on a random seed, encouraging the LLM to choose tokens from the green list. Then, users who know the partition mode can implement the verification by calculating the number of green tokens in the generated text. However, Hu et al. (Hu et al., 2023) found a trade-off between watermark strength and output perplexity. Thus, they embedded watermarks by manipulating the probability distribution of the next token, preferring specific tokens without changing the text semantics. Similarly, Liu et al. (Liu et al., 2023e) proposed a semantic-invariant watermark scheme, generating watermark logits based on the semantics of all preceding tokens instead of simply token IDs. It maintains the semantics of generated texts, successfully resisting synonym replacement attacks and text paraphrasing attacks. Additionally, they found that these watermark schemes used the same key for generation and verification, which posed a forgery risk in the public verification scenario. Inspired by this observation, literature (Liu et al., 2023d) improved the watermarking scheme proposed by Kirchenbauer et al. (Kirchenbauer et al., 2023). They adopted two neural networks for watermark generation and verification, improving the accuracy of verification and the non-forgery of watermarks. Kuditipudi et al. (Kuditipudi et al., 2023) found that this scheme caused the LLM to generate identical watermarked texts for the same prompt, affecting the LLM’s utility. They used longer pseudorandom number sequences to solve this issue and randomly set the initial position during watermark embedding. Subsequently, Hou et al. (Hou et al., 2023) focused on sentence-level sampling and injected watermarks into it, making the watermark resistant to text paraphrasing attacks.
9.3. Machine unlearning for LLMs.
As users increasingly focus on personal privacy protection, especially under legal frameworks such as the GDPR, the right to forget has become a critical problem. Upon receiving a forgetting request, the model owner can leverage machine unlearning to protect their data from unauthorized use. This technology effectively removes the influence of specific user data. Machine unlearning for LLMs faces three challenges compared to traditional small-scale language models. First, LLMs are usually trained on a large amount of data. However, the contribution of a single sample point is not significant, making it challenging to exact unlearning for specific data. Second, machine unlearning will affect the performance of LLMs, requiring a balance between protecting user privacy and maintaining model performance. Third, concerning LLMs, there needs to be more research on verifying the effectiveness of forgetting. Currently, Yao et al. (Yao et al., 2023) studied machine unlearning for LLMs. They found that the memory capabilities of LLMs far exceeded those of traditional models, running multiple unlearning operations to eliminate the impact of specific data. They also identified the issue of catastrophic forgetting, which affects model performance. Meanwhile, Eldan et al. (Eldan and Russinovich, 2023) addressed copyright issues in corpora, replacing ‘Harry Potter’ with other concepts. They made the target LLM forget content related to ‘Harry Potter’. In terms of verification, Viswanath et al. (Viswanath et al., 2024) evaluated the model’s response to forgot samples. In addition, they explored some verification methods, such as data extraction attacks. Despite the many challenges machine unlearning and verification face, research in this area is crucial for improving the transparency of LLMs. Many countries have introduced regulations on AI models, aiming to improve the transparency of AI-based applications. With transparency, it is easier for researchers to understand the inference of LLMs. Improving transparency will be conducive to further research on LLMs’ privacy and security. Mitchell et al. (Mitchell et al., 2019) improve the transparency of models by establishing model cards. Then, Wahle et al. (Wahle et al., 2023) proved that model cards were suitable for LLMs. It helps users better understand model performance and potential risks to train and deploy LLMs more safely.
10. Conclusion
This paper provides a comprehensive overview of privacy risks, security risks, and defenses, focusing on LLMs. Specifically, we introduce various privacy and security attacks, including training-time attacks and inference-time attacks, along with their respective subcategories. We also discuss general and specific protection strategies for privacy and security risks. Furthermore, we present LLMs’ evaluation and provide details on the datasets and evaluation metrics regarding utility and robustness. Concerning LLM privacy and security aspects, this paper gives a comprehensive view of attacks and defenses; we believe it is a significant contribution to help researchers understand LLMs’ issues, making them be applied to more fields.